# Learn the Ropes, Then Trust the Wins: Self-imitation with Progressive Exploration for Agentic Reinforcement Learning

### v2

@@ Abstract
Reinforcement learning (RL) is the dominant paradigm for sharpening strategic tool use capabilities of LLMs on long-horizon, sparsely-rewarded agent tasks, yet it faces a fundamental challenge of exploration-exploitation trade-off. Existing studies stimulate exploration through the lens of policy entropy, but such mechanical entropy maximization is prone to RL training instability due to the multi-turn distribution shifting. In this paper, we target the progressive exploration-exploitation balance under the guidance of the agent’s own experiences without succumbing to either entropy collapsing or runaway divergence. We propose SPEAR [Uncaptioned image], a curriculum-based self-imitation learning (SIL) recipe for training agentic LLMs. It extends the vanilla SIL framework, where a replay buffer stores self-generated promising trajectories for off-policy update, by gradually steering the policy evolution within a well-balanced range of entropy across stages. Specifically, our approach incorporates a curriculum to manage the exploration process, utilizing intrinsic rewards to foster skill-level exploration and facilitating action-level exploration through SIL. At first, the auxiliary tool call reward plays a critical role in the accumulation of tool-use skills, enabling broad exposure to the unfamiliar distributions of the environment feedback with an upward entropy trend. As training progresses, self-imitation gets strengthened to exploit existing successful patterns from replayed experiences for comparative action-level exploration, accelerating solution iteration without unbounded entropy growth. To further stabilize training, we recalibrate the advantages of experiences in the replay buffer to address the potential policy drift. Regularizations such as the clipping of tokens with high covariance between probability and advantage are introduced to the trajectory-level entropy control to curb over-confidence. We also combine bag-of-tricks, which are effective industrial optimizations of agentic RL, into a strong baseline, Dr.BoT, to demonstrate the practical effectiveness of our proposed SPEAR. In ALFWorld and WebShop, SPEAR increases the success rates of GRPO/GiGPO/Dr.BoT by up to 16.1%/5.1%/8.6% and 20.7%/11.8%/13.9%, respectively. In AIME24 and AIME25, SPEAR boosts Dr.BoT by up to 3.8% and 6.1%, respectively. Such gains incur only 10%–25% extra theoretical complexity and negligible runtime overhead in practice, demonstrating the plug-and-play scalability of SPEAR.
@@ 1 Introduction
Reinforcement Learning (RL) [1, 2, 3] has driven the development of reasoning capabilities of Large Language Models (LLMs). Built upon the reason-and-act (ReAct) paradigm [4], LLMs have powered various agentic applications such as simulated robot navigation [5, 6], mobile assistant [7, 8], web navigator [9, 10], deep searcher [11, 12, 13], and GUI master [14, 15]. A fundamental challenge in applying RL to LLM agents, especially for long-horizon tasks with sparse rewards, is managing the balance between exploration and exploitation. The LLM agent needs to exploit both its pretrained knowledge and feedback from past interactions to identify and refine strategies that maximize ultimate reward. At the same time, it must explore novel behaviors, using different tools to discover more effective solutions, which involves extending text-based reasoning to tool-augmented reasoning. The interweaving between exploration and exploitation determines the emerging agent’s competence upon convergence.
Figure 1: The core concept of our proposed SPEAR for training long-horizon LLM agents via group-based RL. Compared with the vanilla GRPO-like algorithms, we introduce the curriculum-based self-imitation learning with intrinsic reward shaping. Given the same data input, a group of trajectories are generated with multi-turn tool interactions and then engaged for episode-level reward computation and advantage estimation. Then, we propose filtering valuable good trajectories to update the replay buffer, where the stored past experiences guide the agent to explore effectively on sparsely rewarded tasks via self-imitation. The total training batch contains both on-policy and off-policy data from the replay buffer. With self-guided progressive exploration, SPEAR boosts performance of various LLMs and baselines on ALFWorld, WebShop, and AIME24/25 in a plug-and-play manner.
@@
Existing studies often quantify the exploration potential through entropy [16, 17, 18], where the decline of policy entropy indicates the over-confidence of the current policy with insufficient exploration. In this case, a series of regularization techniques [19, 20, 21] have been proposed to maximize entropy, particularly in the context of agent training [22, 23, 24, 25, 26, 27, 28, 29]. However, when it comes to LLM-driven agents, entropy-based control is fragile: the accumulation of low-probability tokens from the environment feedback induces severe distribution shifting, often leading to mode collapse [18, 30]. Unlike conventional agents, these models may experience sustained entropy growth due to policy over-uncertainty during multi-turn interactions with external tools. As a result, the policy struggles to exploit tool-use skills effectively, and training instability becomes frequent [31, 32, 33]. Recent approaches attempt to mitigate this issue by relying either on cold-start supervised fine-tuning (SFT) [13, 14, 34, 35] or on hybrid schemes that combine RL with SFT [36]. Although these methods improve stability, they compromise RL’s ability to discover strategies beyond those present in the SFT corpus. This limitation highlights the need for adaptive training frameworks that can dynamically schedule LLM-driven agents to decide when to explore and when to exploit.
In this paper, we are trying to answer the following core research question:
During RL of agentic LLMs, can we schedule a smooth transition between exploration and exploitation guided by the policy’s own experiences without going to extremes of either entropy collapsing or runaway divergence?
We hypothesize that policy entropy should serve as a progressive guide for balancing exploration and exploitation throughout training. Specifically, the agent should maintain its entropy within a dynamic but controlled range that evolves over time:
  • 1) At the early stages of training, increasing entropy is beneficial for supporting broad skill-level exploration and rapidly developing tool-use capabilities. The agent is expected to encounter unfamiliar observations, engage in trial-and-error interactions, and discover diverse reasoning trajectories.
  • 2) As training advances, however, a shift toward converging entropy is required. This enables the agent to consolidate problem-solving heuristics and emphasize action-level exploration. In this phase, the agent exploits reward signals to choose comparatively more effective actions and adapts to changing trajectory distributions, thereby stabilizing its evolutionary path.
@@
Figure 2: Overview of SPEAR in terms of data flow. During each episode, the agent interacts with the environment to generate a set of trajectories. These trajectories are processed along two complementary paths. First, they are used for intrinsic reward shaping, advantage estimation, and on-policy updates, following a mechanism similar to the vanilla GRPO. Second, they are selectively filtered and stored in a replay buffer, enabling off-policy updates through the proposed self-imitation scheme with advantage recalibration and regularization. This dual integration allows the agent to maximize the utility of rewarding past experiences, thereby expanding the exploration space effectively, while simultaneously mitigating persistent over-uncertainty in decision-making under shifting distributions of external feedback. As a result, SPEAR achieves a stable balance between exploration and exploitation through self-guided policy adaptation.
@@
To address this, we propose the Self-imitation with Progressive Exploration for Agentic Reinforcement learning (SPEAR)[Uncaptioned image], a curriculum-based RL recipe for improving the exploration-exploitation balance with self-imitation and intrinsic reward. As shown in Figure 1, the core principle of philosophy follows the vanilla Self-Imitation Learning (SIL) [37, 38] where an independent replay buffer is prepared to store the state-action pairs (i.e., trajectories) only when their returns in the past episodes exceed the baselines (i.e., those with positive advantages). Such a replay buffer is exploited to encourage actions with good returns and improve hard exploration based on these successful trajectories under the sparse-reward, long-horizon agent tasks. Specifically, we introduce three modifications to SIL tailored to the dynamics of policy entropy in agentic tasks. First, we incorporate a curriculum to integrate both skill-level and action-level exploration by adjusting reward shaping and self-imitation across stages. For skill-level exploration, intrinsic rewards encourage frequent tool usage, pushing the agent to interact with uncertain environments. At early stages, exploration remains unconstrained to avoid overfitting immature experiences in the replay buffer. For action-level exploration, we progressively amplify self-imitation of promising trajectories, allowing the agent to iteratively refine behaviors while preventing unbounded entropy growth. To ensure outcome reward remains dominant, we gradually reduce the weight of intrinsic rewards. Second, we tackle the off-policy nature of the update with experiences in the buffer and avoid advantage recomputation by advantage recalibration. Unlike GRPO-like methods [39, 40] that estimate baselines via intra-group averages or the vanilla SIL algorithm [37, 38] that require re-estimating the advantage, we compensate for performance drift in the replay buffer under the assumption of iterative policy improvement to avoid computation. Third, we regularize policy updates to stabilize entropy and mitigate reward hacking. Specifically, we introduce covariance-based clipping [17] at the trajectory level, where high-probability actions strongly correlated with advantage gains are excluded from gradient contributions. This prevents over-confidence in narrow success patterns while avoiding excessive uncertainty. Finally, inspired by existing industrial “bag-of-tricks,” in LLM-based agentic RL, we present a strong baseline, Dr.BoT. It is yet another GRPO variant combining stability- and efficiency-oriented techniques. Empirical results demonstrate its effectiveness across diverse LLM architectures and agentic tasks.
We show that our proposed SPEAR brings considerable performance gains to multiple baselines, improving over GRPO/GiGPO [41]/Dr.BoT respectively up to 16.1%/5.1%/8.6% on ALFWorld [5] and 20.7%/11.8%/13.9% on WebShop [42]. It boosts our Dr.BoT respectively up to 3.8% on AIME24 and 6.1% on AIME25 [43]. These gains come with around $10\%\sim 25\%$ computation overhead in theoretical complexity, but end up with quite similar and comparable runtime per iteration in practice. Such compatibility and scalability make our SPEAR a plug-and-play algorithm for training versatile LLM agents. In summary, our contributions are:
  • We propose SPEAR, a generalization of the self-imitation learning method for training agentic LLMs. We bring in curriculum scheduling for progressive transition from skill-based to action-based exploration. It bypasses the costly imitation of expert demonstrations and allows the agent to explore for long-term, distal rewards under the guidance of one’s own experiences.
  • We target the management of policy entropy to prevent either premature collapsing or persistent instability. We propose an integration of regularization techniques and intrinsic reward shaping to balance exploration-exploitation via the lens of entropy.
@@
  • We propose a strong GRPO-variant baseline, Dr.BoT, which combines established agentic RL training techniques validated in industrial practice, confirming its effectiveness and superiority over existing baselines.
  • Extensive experiments across LLMs and tasks demonstrate that our SPEAR consistently enhances the performance of various baselines, showcasing its plug-and-play nature with high compatibility and scalability.
@@ 2.1 Reinforcement Learning Algorithms for LLMs
Reinforcement learning (RL) was first introduced into LLM training to improve instruction following and alignment [44]. With the advent of large-scale reasoning models (e.g., OpenAI’s o1 [45]), RL has been adopted more broadly in post-training pipelines. Proximal Policy Optimization (PPO) [20], as the canonical online RL method, leverages an actor–critic architecture together with the clipped surrogate objective and a Kullback–Leibler (KL) divergence penalty to constrain policy update while achieving strong performance. Group Relative Policy Optimization (GRPO) [2, 39] simplifies this setup by replacing the critic with a group-wise baseline: for each prompt, multiple rollouts are sampled to form a group, and their rewards are normalized using the group mean and standard deviation. This critic-free design reduces memory usage and engineering complexity while retaining PPO-like trust-region behavior. Building on GRPO, many variants target the training stability and sample efficiency. DAPO [46] uses dynamic sampling and ”clip higher” to encourage exploration and stabilize training. Dr.GRPO [40] addresses length bias and the difficulty bias by normalization, yielding better token efficiency. GSPO [47] proposes the sequence-level importance ratio to improve training efficiency notably for Mixture-of-Experts (MoE) [48] LLMs.
Existing methods have greatly advanced RL for LLMs. However, naively combining them can lead to conflicts or tight couplings among methods. To this end, we harmonize the strengths of DAPO, Dr.GRPO, and other agent RL studies from both research and industrial practice to establish a strong baseline, Dr.BoT, as detailed in Section 4.3.
@@ 2.2 Optimization of LLM Agents
The success of GRPO for RL has expedited the training of LLM-based agents. A growing line of researches investigate how to endow models with better tool-use capabilities so that they can complete complex tasks [34, 49, 18]. In parallel, recent works [11, 13, 50] connect LLMs to the open web to strengthen information seeking capability with smarter search and browsing. RAGEN [33] improves the stability of multi-turn RL training through instance filtering and gradient shaping. GiGPO [41] augments group-level advantages with additional step-level advantage estimates by identifying shared states across different trajectories, capturing both global and local preference against the sparse reward problem. ARPO [30] monitors entropy dynamics during rollouts and branches trajectories adaptively, promoting exploration at steps with high uncertainty after tool usage. Such branching internalizes advantage differences in stepwise tool-calls like GiGPO.
In this work, we address the exploration–exploitation dilemma under multi-turn tool-use settings. We introduce a curriculum–regulated RL regime that gradually shifts skill-based exploration towards action-based exploration. We integrate self-imitation and intrinsic reward to consolidate successful behaviors, which will be discussed in Section 4.1. It is noted that our SPEAR can work with existing algorithms in a plug-and-play manner, exhibiting a high level of compatibility and generalization for policy entropy control.
@@ 2.3 Exploration in Reinforcement Learning
In sparse or delayed environments, it is pivotal to motivate the agents to explore unknown states for sample efficiency and generalization. Curiosity-driven method [51] grant intrinsic rewards for prediction error or novelty to actively seek unfamiliar states without extrinsic feedback. Similarly, VIME [52] rewards actions that maximize the information gains about the agent’s belief of the environment’s dynamics, improving exploration in continuous control task. Count-based algorithms encourage exploration by awarding higher intrinsic rewards for visiting rare or previously unseen states. [53] introduces pseudo-counts derived from a density model to assign count-based bonuses in high-dimensional problems. [54] shows that a simple hashing of discrete states for counting enables strong exploration in continuous control and visual games. Skill acquisition methods learn a diverse repertoire of skills or options through intrinsic objectives, so that these behaviors can later aid exploration or downstream tasks. VIC [55] discovers a set of distinct options by maximizing the mutual information and can effectively learn skills that reliably reach different regions of the state space. DIAYN [56] learns useful skills without any extrinsic reward by maximizing an information-theoretic objective under a maximum-entropy policy. Entropy-regularization methods explore high-entropy actions to prevent premature convergence. SAC [21] establishes a framework where the agent explicitly maximizes the expected reward and entropy. [17] proposes to control entropy by restricting the update of high-covariance tokens to prevent entropy collapsing of LLMs.
However, in the era of agentic LLMs, the naive application of entropy maximization could lead to divergence as the multi-turn interactions already result in the increased uncertainty on unfamiliar observations. Under such circumstance, we propose the curriculum-guided self-imitation to leverage the agent’s own experiences for balancing exploration and exploitation, where the intrinsic reward and self-imitation respectively enhance skill-based and action-based exploration. It avoids handcrafted heuristic techniques in previous studies and instead fully relies on the agent itself to reinforce successful patterns for efficiency of valid exploration.
@@ 2.4 Experience Replay in Reinforcement Learning
Self-Imitation Learning (SIL) [37] aims to take advantage of the past successful experiences [57, 58, 59, 60] to drive its future learning. The core insight is that learning from one’s own high-return trajectories could indirectly encourage exploration in environments with sparse rewards. Building upon SIL, [61] formulates policy optimization as a divergence minimization problem, which can be further reduced into a policy-gradient algorithm with shaped rewards learned from experience replays. Self-imitation advantage learning (SAIL) [38] extends SIL to off-policy, action-value-based RL methods. [62] proves that SIL’s return-based update provides a bias–variance trade-off that can speed up learning by propagating rewards over longer horizons. Self-imitation learning from demonstrations (SILfD) [63] extends SIL to leverage external demonstrations and the agent’s own experiences for the improved performance. For training LLMs, Generalized SIL (GSIL) [64] proposes an offline alignment framework that uses self-imitation on demonstration data. By defining a surrogate loss with density ration estimation, GSIL allows the model to learn from its own generated solutions alongside reference answers. The success of SIL in both game environments and modern LLM applications shows that agents can effectively learn from themselves by leveraging their own successful experiences as guideposts for future behavior.
While SIL helps on long-horizon problems, its direct application on agentic RL induces entropy collapsing. To mitigate this, we introduce an intrinsic reward to sustain skill-based exploration and harmonize its contribution with self-imitation via curriculum. We additionally suppress over-confidence by clipping high covariance tokens out of SIL as regularization.

### v3

@@ Abstract
Reinforcement learning (RL) is the dominant paradigm for sharpening strategic tool use capabilities of LLMs on long-horizon, sparsely-rewarded agent tasks, yet it faces a fundamental challenge of exploration-exploitation trade-off. Existing studies stimulate exploration through the lens of policy entropy, but such mechanical entropy maximization is prone to RL instability due to the multi-turn distribution shifting. In this paper, we target the progressive exploration-exploitation balance under the guidance of the agent’s own experiences without succumbing to either entropy collapsing or runaway divergence. We propose SPEAR [Uncaptioned image], a self-imitation learning (SIL) recipe for training agentic LLMs. It extends the vanilla SIL, where a replay buffer stores good experience for off-policy update, by gradually steering the policy entropy across stages. Specifically, the proposed curriculum scheduling harmonizes intrinsic reward shaping and self-imitation to 1) expedite exploration via frequent tool interactions at the beginning, and 2) strengthen exploitation of successful tactics upon convergence towards familiarity with the environment. We also combine bag-of-tricks of industrial RL optimizations for a strong baseline Dr.BoT to demonstrate our effectiveness. In ALFWorld and WebShop, SPEAR increases the success rates of GRPO/GiGPO/Dr.BoT by up to 16.1%/5.1%/8.6% and 20.7%/11.8%/13.9%, respectively. In AIME24 and AIME25, SPEAR boosts Dr.BoT by up to 3.8% and 6.1%, respectively. Such gains incur only 10%–25% extra theoretical complexity and negligible runtime overhead in practice, demonstrating the plug-and-play scalability of SPEAR.
@@ 1 Introduction
Figure 1: Our SPEAR harmonizes the curriculum-scheduled self-imitation learning with intrinsic reward shaping for progressive exploration, improving policy performance across agentic tasks.
@@
Reinforcement Learning (RL) [1, 2, 3] has driven the development of reasoning capabilities of Large Language Models (LLMs). Built upon the reason-and-act (ReAct) paradigm [4], LLMs have powered various agentic applications such as simulated robot navigation [5, 6], mobile assistant [7, 8], web navigator [9, 10], deep searcher [11, 12, 13], and GUI master [14, 15]. A fundamental challenge in applying RL to LLM agents is to manage the balance between exploration and exploitation. The LLM agent needs to exploit both pretrained knowledge and past interactions to formalize experience that maximize rewards. At the same time, it must explore novel behaviors through tool-integrated reasoning and reflection. The interweaving between exploration and exploitation determines the emerging agent’s competence upon convergence.
@@
Existing studies often quantify the exploration potential through entropy [16, 17, 18, 19], where the decline of policy entropy indicates over-confidence with insufficient exploration. In this case, a series of regularization techniques [20, 21, 22] have been proposed to maximize entropy  [23, 24, 25, 26, 27, 28, 29, 30]. However, when it comes to LLM-driven agents, entropy-based control is fragile: the accumulation of low-probability tokens from the environment feedback induces severe distribution shifting, often leading to mode collapse [19, 31]. Agent models may experience sustained entropy growth due to uncertainty about multi-turn interactions and training instability becomes frequent [32, 33, 34]. Recent approaches attempt to mitigate this issue by cold-start supervised fine-tuning (SFT) [13, 14, 35, 36] or hybrid schemes that combine RL with SFT [37]. Although these methods improve stability, they compromise policy’s discovery of strategies beyond those present in the SFT corpus. This limitation highlights the need for adaptive training frameworks that can dynamically schedule LLM-driven agents to decide when to explore and when to exploit.
In this paper, we are trying to answer the following core research question: Can we schedule a smooth transition between exploration and exploitation guided by the policy’s own experience without going to extremes of either entropy collapsing or runaway divergence? We hypothesize that the agent should maintain its policy entropy within a dynamic but controlled range that evolves over time: 1) At the early stages, increasing entropy is beneficial for broad skill-level exploration. The agent is expected to rapidly develop tool-use capabilities, encounter unfamiliar observations, and engage in trial-and-errors. 2) As training advances, however, a shift toward converging entropy is required. This enables the agent to consolidate problem-solving heuristics and emphasize action-level exploration. The agent exploits reward signals to choose comparatively more effective actions and adapts to changing distributions for stabilizing its evolutionary path.
@@
Figure 2: Overview of SPEAR. First, the agent interacts with the environment for a set of trajectories, which flow through intrinsic reward shaping and advantage estimation with on-policy updates. Second, they are selected and stored in a replay buffer, enabling off-policy updates via the proposed self-imitation scheme. This dual integration allows the maximal utility of past experiences, thereby expanding the effective exploration space, while simultaneously mitigating persistent uncertainty.
@@
To address this, we propose the Self-imitation with Progressive Exploration for Agentic Reinforcement learning (SPEAR[Uncaptioned image]), a curriculum-based RL recipe for improving the exploration-exploitation balance with self-imitation and intrinsic reward. As shown in Figure 1, the core principle follows the vanilla Self-Imitation Learning (SIL) [38, 39] where an independent replay buffer is prepared to store the state-action pairs only when their returns in the past episodes exceed the baselines. Such a replay buffer is exploited to encourage actions with good returns and improve hard exploration based on these successful trajectories under the sparse-reward, long-horizon agent tasks. Specifically, we introduce three modifications to SIL tailored to the dynamics of policy entropy in agentic tasks. First, we incorporate a curriculum to integrate both skill-level and action-level exploration by adjusting reward shaping and self-imitation across stages. Second, we tackle the off-policy nature of the update with experiences in the buffer and avoid advantage recomputation by advantage recalibration. Third, we regularize policy updates to stabilize entropy and mitigate reward hacking. Finally, inspired by existing industrial bag-of-tricks, we present a strong baseline Dr.BoT for agentic RL training. Our SPEAR brings considerable performance gains to GRPO/GiGPO [40]/Dr.BoT respectively up to 16.1%/5.1%/8.6% on ALFWorld [5] and 20.7%/11.8%/13.9% on WebShop [41]. It boosts our Dr.BoT respectively up to 3.8% on AIME24 and 6.1% on AIME25 [42]. These gains come with around $10\%\sim 25\%$ computation overhead in theoretical complexity, but end up with quite comparable runtime per iteration in practice. Such compatibility and scalability enable SPEAR a plug-and-play algorithm for training versatile agents. In summary, our contributions are:
1) We propose SPEAR, a generalization of the SIL for training LLM agents. It bypasses the costly expert imitation and allows exploration under the guidance of one’s own rewarded experience.
2) We bring in curriculum scheduling to harmonize SIL with intrinsic reward shaping for policy entropy management and progressive transition from skill-based to action-based exploration.
@@
3) We propose a strong baseline, Dr.BoT, which combines established RL techniques validated in industrial practice, confirming its effectiveness and superiority over existing baselines.
@@ 2.1 Reinforcement Learning Algorithms for LLMs
With the advent of large-scale reasoning models [43], Reinforcement Learning (RL) [44] has been adopted more broadly. Proximal Policy Optimization (PPO) [21] leverages an actor–critic architecture together with the clipped surrogate objective and a Kullback–Leibler (KL) divergence penalty to constrain policy update. Group Relative Policy Optimization (GRPO) [2, 45] simplifies this setup by replacing the critic with a group-wise baseline. Building on GRPO, DAPO [46] uses dynamic sampling and ”clip higher” to encourage exploration and stabilize training. Dr.GRPO [47] addresses length bias and the difficulty bias. Existing methods have greatly advanced RL for LLMs. However, naively combining them can lead to conflicts or tight couplings among techniques. To this end, we harmonize the strengths of DAPO, Dr.GRPO, and other agent studies from research and industrial practice to establish a strong baseline, Dr.BoT, as detailed in Section 4.4.
@@ 2.2 Optimization of LLM Agents
Recent researches investigate how to endow models with better tool-use capabilities [35, 48, 19]. LLMs are optimized to strengthen information seeking from open web [11, 13, 49]. RAGEN [34] improves the stability of multi-turn RL through instance filtering and gradient shaping. GiGPO [40] augments group-level advantages with additional step-level advantage estimates. ARPO [31] monitors entropy dynamics during rollouts to branch trajectories adaptively. In this work, we address the exploration–exploitation dilemma under multi-turn tool-use settings. We introduce a curriculum–regulated RL regime that gradually shifts skill-based exploration towards action-based exploration. We integrate self-imitation and intrinsic reward to consolidate successful behaviors (Section 4.2). Our SPEAR can work with existing algorithms in a plug-and-play manner, exhibiting a high level of compatibility and generalization.
@@ 2.3 Exploration in Reinforcement Learning
Curiosity-driven methods [50, 51] grant intrinsic rewards for prediction error or novelty to actively seek unfamiliar states. Count-based algorithms [52, 53] introduce pseudo-counts derived from a density model to assign count-based bonuses. Skill acquisition methods [54, 55] discover distinct options by maximizing the mutual information. Entropy-regularization methods [22, 18] maximize the expected reward and entropy of the policy. However, traditional exploration techniques can lead to divergence of agent LLMs as the multi-turn interactions already result in the increased uncertainty on unfamiliar observations. Under such circumstance, we propose the curriculum-guided self-imitation to leverage the agent’s own experience for balancing exploration and exploitation. It avoids handcrafted heuristic techniques in previous studies and instead fully relies on the agent itself to reinforce successful and valid patterns.
@@ 2.4 Experience Replay in Reinforcement Learning
Self-Imitation Learning (SIL) [38] takes advantage of past successful experience to drive its future learning [56, 57, 58, 59, 60]. SAIL [39] extends SIL to off-policy, action value-based RL methods. [61] proves that SIL’s return-based update provides a bias–variance trade-off that speeds up learning. SILfD [62] extends SIL to leverage both external demonstrations and the agent’s own experience. GSIL [63] proposes an offline alignment framework that uses self-imitation on demonstration data. While SIL benefits long-horizon problems, its induces entropy collapsing to agent RL. To mitigate this, we harmonize both self-imitation and intrinsic reward with curriculum scheduling for progressive exploration.